OpenStudy (anonymous):

how do you do n -8= -2? - 3

OpenStudy (anonymous):

i lied 3n-8=-6 add the 8 3n=2 divide by 3 n=2/3

OpenStudy (anonymous):

its actually n over 3 minus 8= -2...

OpenStudy (anonymous):

okay. easy fix. n/3=6 times by 3. n=18

OpenStudy (anonymous):

thank you so much!